DX on Sky Checkout

02:37
  • Waiting for Arcetri group to finish Cal-Unit work.
03:31
  • Telescope now open, Steve is checking on something.
05:08
  • Lost most of the log due to an issue with X. (I opened the "Big Display" and couldn't close it or get my desktop back)
  • It mainly had to do with Doug working on preset issues.
05:23
  • Bin 4 15.0 mag Double!
  • Data with 10 LUCI frames.
  • 052613, 052748, 052913, 053028 Loop opened.
05:29
  • Get rid of excess ccd39 viewers with: ps aux | perl -e 'while(<>){if(/^[A-Za-z]+\s+([0-9]+).+ccd_viewer\.py\s+ccd39$/){print "$1\n"}}' | xargs kill -9
05:38
  • New preset, Bin3, 13.2
  • Datasets: 055259, 055330, 055353, 055438
  • Datasets: with 1 sec. 060024 Bad, 060117, 060145, 060208, 060303
06:07
  • Starting a dice 5
  • Had a safeskip open the loop on one of the offsets.
  • Was able to use "Re-close Loop"
06:37
  • Bin 2 11.8 mag
  • GOPT on 800 Hz
06:55
  • New preset, same star.
  • Doing Dice 5
07:21
  • Reclose loop + optimize gain on the FLAO gui resets the optical gain to 1.0, but it doesn't turn NCPA off.
  • This will cause the loop to open with "Safe Skip" because the NCPA amplitude applied will be 2-3X too much.
  • Actuator 514 made the shell rip after several jumps tonight, removed it in position. (added to no position list)
07:50
  • Re-doing preset manually.
08:14
  • Closing due to humidity

-- %USERSIG{GregoryTaylor - 2017-09-21}%

Comment

No Comment
Topic revision: r4 - 21 Sep 2017, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback